OpenStudy (anonymous):

can someone help me with the limit: lim as x approaches2 of: (sqrt x^2-4x+4)/(x-2)

OpenStudy (anonymous):

sqrt(x-2)^2/(x-2)=1

OpenStudy (anonymous):

could soemone type it out for me, thanks.

OpenStudy (anonymous):

Factor the top giving: \[\lim_{x \rightarrow 2} \frac{\sqrt{(x-2)^2}}{x-2}=\lim_{x \rightarrow 2}1=1\]

OpenStudy (anonymous):

dont we get some absolute value type thing, like the (x-2) is actually + or - (x-2)

OpenStudy (anonymous):

Only if you take a sqrt, but since it is already there you use the sign convention in the problem. If it was +/- then the limit MIGHT not exist. (I.e., it would approach two different values)

OpenStudy (anonymous):

this limit be undefined

OpenStudy (anonymous):

this limit be undefined

OpenStudy (anonymous):

you are taking \[\lim_{x->2}\frac{|x-2|}{x-2}\]

OpenStudy (anonymous):

how is it undefined?

OpenStudy (anonymous):

because your function is \[ f(x) = \left\{ \begin{array}{lr} 1 & : x >2\\ -1 & : x <2 \end{array} \right. \] and the limit at 2 does not exist since the right hand limit is not equal to the left hand limit

OpenStudy (anonymous):

see, i knew it

OpenStudy (anonymous):

oh yes. i see you stated that above. good work.

OpenStudy (anonymous):

when we take sqrt of (x-2)^2, we end up with absolute value of (x-2), which could either be + or -(x-20

OpenStudy (anonymous):

right. the function is exactly what you said, which translates into the piece-wise function that i wrote above.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

could you help me with the other question i posted regarding simplifiying?

OpenStudy (anonymous):

it is just like the usual example of \[\frac{|x|}{x}\] which has no limit at 0 yw

OpenStudy (anonymous):

back in ten minutes then i will look

OpenStudy (anonymous):

k